OpenStudy (anonymous):

Find the area of a circle with a diameter of 24 inches and use 3.14 for pi

OpenStudy (lgbasallote):

area = pi (d/2)^2 area = (3.14)(24/2)^2 area = (3.14)(12)^2 does that help?

OpenStudy (anonymous):

A=pi . r^2 A=3.14(12) multiply 3.14 and 12 the result is the answer

OpenStudy (lgbasallote):

you have to square 12 @cornitodisc

OpenStudy (anonymous):

Thank you so much. Yes, that helps. I thought that's what I had to do, but got confused when I saw something else on the internet that they did not square

OpenStudy (lgbasallote):

you must have seen circumference